# Learning to Evolve: Bayesian-Guided Continual Knowledge Graph Embedding

### v1

@@ Introduction
Inspired by related work such as uncertainty regularized(Li et al. 2025a) continual learning(Ahn et al. 2019)(Lee et al. 2024)(Bonnet et al. 2025), we regard the embedding parameters of entities and relations as random variables in probability distributions rather than fixed point estimates. This approach not only facilitates quantification of parameter uncertainty but also protects important old knowledge when learning new knowledge, avoiding catastrophic forgetting.
@@
Entity and Relation Distributional Representations For each entity $e$ and relation $r$, their embedding vectors at time snapshot $t$ are modeled as Gaussian distributions. The embedding of entity $e$ is represented as $\mathbf{e}_{t}\sim\mathcal{N}(\boldsymbol{\mu}_{e,t},(\text{diag}(\boldsymbol{\lambda}_{e,t}))^{-1})$, where $\boldsymbol{\mu}_{e,t}\in\mathbb{R}^{d}$ is the mean vector, and $\boldsymbol{\lambda}_{e,t}\in\mathbb{R}^{d}$ is the diagonal precision vector (inverse of variance); the embedding of relation $r$ is similar, represented as $\mathbf{r}_{t}\sim\mathcal{N}(\boldsymbol{\mu}_{r,t},(\text{diag}(\boldsymbol{\lambda}_{r,t}))^{-1})$. Each element of the precision $\boldsymbol{\lambda}$ reflects the certainty of the corresponding dimension: a higher value indicates more reliable knowledge in that dimension, which should be better protected in subsequent updates.
@@
Sequential Bayesian Update At time snapshot $t$, we use the posterior distribution $\{\boldsymbol{\mu}_{t-1},\boldsymbol{\lambda}_{t-1}\}$ from time $t-1$ as the prior and perform online Bayesian update using new data $\Delta\mathcal{T}_{t}$. To simplify inference, the embeddings $\hat{\boldsymbol{\Theta}}_{t}$ obtained from training on the current snapshot are treated as new observations to the prior mean $\boldsymbol{\mu}_{t-1}$, generated by the classic KGE model TransE (Bordes et al. 2013). TransE evaluates the plausibility of triples through a scoring function and optimizes the embeddings $\hat{\boldsymbol{\Theta}}_{t}=\{\hat{\mathbf{e}}_{t},\,\hat{\mathbf{r}}_{t}\}$ by minimizing the margin-based ranking loss $\mathcal{L}_{\text{KGE}}$:
$\mathcal{L}_{\mathrm{KGE}}=\sum_{(h,r,t)\in\mathcal{T}}\sum_{\left(h^{\prime},r,t^{\prime}\right)\in\mathcal{T}^{-}}\left[\gamma+f(h,r,t)-f\left(h^{\prime},r,t^{\prime}\right)\right],$ (1)
@@
where $f(h,r,t)=\lVert\mathbf{h}+\mathbf{r}-\mathbf{t}\rVert_{2}$ represents the score function of TransE(Bordes et al. 2013).Based on the conjugate property of Gaussian distributions, the update rules for the embedding parameters of entity $e$ are:
$\boldsymbol{\lambda}_{e,t}$ $=\boldsymbol{\lambda}_{e,t-1}+\lambda_{obs},$ (2)
$\boldsymbol{\mu}_{e,t}$ $=\frac{\boldsymbol{\lambda}_{e,t-1}\odot\boldsymbol{\mu}_{e,t-1}+\lambda_{obs}\odot\hat{\mathbf{e}}_{t}}{\boldsymbol{\lambda}_{e,t}},$ (3)
@@
To guide the embeddings towards the posterior mean, we use a regularization term (Lee et al. 2024) $\mathcal{L}_{Bayes}$, which is equivalent to minimizing the KL divergence between the current distribution $q(\mathbf{\Theta}_{t})$ and the target posterior:
$\mathcal{L}_{Bayes}=\sum_{i\in\mathcal{E}_{t}\cup\mathcal{R}_{t}}\beta\cdot\left\|\sqrt{\boldsymbol{\lambda}_{i,t-1}}\odot(\hat{\boldsymbol{\theta}}_{i,t}-\boldsymbol{\mu}_{i,t-1})\right\|_{2}^{2},$ (4)
@@
Snapshot 0 Snapshot 1 Snapshot 2 Snapshot 3 Snapshot 4
Dataset $N_{E}$ $N_{R}$ $N_{T}$ $N_{E}$ $N_{R}$ $N_{T}$ $N_{E}$ $N_{R}$ $N_{T}$ $N_{E}$ $N_{R}$ $N_{T}$ $N_{E}$ $N_{R}$ $N_{T}$
Table 1: The statistics of datasets. $N_{E}$, $N_{R}$ and $N_{T}$ denote the number of entities and relations, and the current snapshot $i$
Continual Clustering
@@
Bayesian updates prevent forgetting at the parameter level, but the evolution of knowledge representations between KG snapshots needs to be constrained; otherwise, entity and relation embeddings in the latent space may undergo uncertainty drift across snapshots. Inspired by contrastive learning and clustering methods (Liang et al. 2024a)(Truong et al. 2025)(Nguyen et al. 2021), we propose a continual clustering method as a constraint to maintain the geometric structural consistency of the embedding space at the semantic level.
@@
Sequential Contrastive Clustering Learning In continual learning, imbalances in data volume and class distribution across temporal snapshots may lead to under-learning of minority classes. To address this, we propose the sequential contrastive clustering method ($\mathcal{L}_{FCC}$), which achieves fairness by dynamically adjusting the contributions of major and minor classes. Specifically, for each cluster $k$, we maintain a centroid vector $\mathbf{c}_{k}\in\mathbb{R}^{d}$ (where $d$ is the embedding dimension) to represent the central features of the cluster. First, we acknowledge the perspective proposed in (Liu et al. 2024a) that the learning order is also crucial in the process of learning KG structures. Second, since new entities emerge in each snapshot, making it difficult to apply k-means clustering, we sort the entities based on their importance, then fix the size of each cluster $k$, and take the mean of the embeddings in the current cluster $k$ as the clustering centroid $\mathbf{c}_{k}$. Specifically, the sorting process is as follows: we comprehensively consider the node’s centrality and betweenness centrality in the graph structure to compute its importance score. The calculation formulas for these two metrics are as follows:
$f_{nc}(e)=\frac{f_{\text{neighbor}}(e)}{N-1},\quad f_{bc}(e)=\sum_{s,t\in\mathcal{E},s\neq t}\frac{\sigma(s,t|e)}{\sigma(s,t)}$ (5)
@@
where $f_{nc}(e)$ reflects the tightness of local connections for the entity, and $f_{bc}(e)$ measures its bridging role in information propagation. The variables in the formulas are defined as follows: $f_{\text{neighbor}}(e)$ represents the number of neighbors of entity $e$, $N$ is the total number of entities in the current knowledge graph snapshot, $\mathcal{E}$ is the set of entities in the current time snapshot $i$, $\sigma(s,t)$ is the total number of shortest paths between entities $s$ and $t$, and $\sigma(s,t\mid e)$ is the number of those paths that pass through entity $e$. We combine the two to define the importance score for entity $e$ as: $IE(e)=f_{nc}(e)+f_{bc}(e)$. We calculate the IE score for each entity and sort them in descending order of scores. Then, we assign the sorted entities to clusters of fixed size in sequence, and compute the initial embedding for each cluster as $\mathbf{c}_{k}=\frac{1}{|K|}\sum_{\mathbf{e}_{i}\in K}\mathbf{e}_{i}$. The continual clustering then constrains the representations of entities across different snapshots through contrastive loss, with the loss function defined as:
@@
$\mathcal{L}_{\text{FCC}}=-\sum_{k=1}^{K}\Bigl{(}\alpha_{k}\!\!\sum_{\mathbf{e}_{i}\in\text{Cluster}_{k}}\mathcal{L}_{\text{cont}}(\mathbf{e}_{i},\mathbf{c}_{k})+\mathcal{L}_{\text{cont}}(\mathbf{v}_{k},\mathbf{c}_{k})\Bigr{)}$ (6)
[…]
@@
$\mathcal{L}_{\text{cont}}\left(\mathbf{e}_{i},\mathbf{c}_{k}\right)=-\log\frac{\exp\left(\operatorname{sim}\left(\mathbf{e}_{i},\mathbf{c}_{k}\right)/\tau\right)}{\sum_{j=1}^{K}\exp\left(\operatorname{sim}\left(\mathbf{e}_{i},\mathbf{c}_{j}\right)/\tau\right)}.$ (7)
@@
Cluster Maintenance and Update At snapshot $t$, the centroids $\mathbf{c}_{k}$ of old classes are inherited and fixed from $t-1$, serving as anchors for old knowledge. New classes initialize centroids via feature means. During training, entities are dynamically assigned to the nearest cluster, and centroids are slowly adjusted through a momentum update:
$\mathbf{c}_{k,t}=(1-\eta)\mathbf{c}_{k,t-1}+\eta\cdot\text{mean}(\{\mathbf{e}_{i}\in\text{Cluster}_{k}\}),$ (8)
@@
Final Training Objective
@@
$\mathcal{L}_{total}=\mathcal{L}_{KGE}+\mathcal{L}_{Bayes}+\mathcal{L}_{FCC}.$ (9)
Experiments
Experimental Setup
Datasets To thoroughly verify the effectiveness and stability of BAKE, we conducted extensive experiments on a wide variety of datasets, including: ENTITY (Cui et al. 2023), RELATION (Cui et al. 2023), HYBRID (Cui et al. 2023), GraphLower (Liu et al. 2024a), GraphEqual (Liu et al. 2024a), GraphHigher (Liu et al. 2024a), FB-CKGE (Liu et al. 2024b), and WN-CKGE (Liu et al. 2024b). These eight datasets encompass different evolution patterns, each with unique characteristics. Detailed statistics of the datasets are listed in Table 1. Each dataset contains 5 snapshots.
Implementation Details All experiments are implemented in PyTorch(Paszke et al. 2019) and conducted on 8 NVIDIA Tesla V100 GPUs. We tune all hyperparameters via grid search, including embedding dimensions for entities and relations 50, 100, 200, 300, batch sizes 256, 512, 1024, and number of clusters $K$ 256, 512, 1024, among others. We use Adam as the optimizer with learning rates selected from 1e-5, 1e-4, 1e-3. To ensure fairness, all reported results are averaged over five runs.
@@
Table 2: Performance comparison of various models on ENTITY, RELATION, FB-CKGE, and WN-CKGE datasets. The best results are in bold, and the second-best results are underlined.
@@
BAKE 0.247 0.147 0.448 0.257 0.154 0.463 0.241 0.142 0.439 0.228 0.146 0.419
Table 3: Performance comparison of various models on GraphLower, GraphEqual, GraphHigher, and HYBRID datasets.
ENTITY RELATION FB‑CKGE WN‑CKGE
@@
Table 4: Ablation study results on ENTITY, RELATION, FB‑CKGE, and WN‑CKGE datasets.
@@
Baselines To fully demonstrate the effectiveness of BAKE, we compare it with all recent CKGE models, including Fin-tune (Cui et al. 2023) (based on fine-tuning with new triples), InDE (Liu et al. 2024a) and FastKGE (Liu et al. 2024b) (based on incremental distillation and LoRA fine-tuning), LKGE (Cui et al. 2023), EWC (Kirkpatrick et al. 2017), SI (Zenke, Poole, and Ganguli 2017), DiCGRL (Kou et al. 2020), EMR (Wang et al. 2019), CWR (Lopez-Paz and Ranzato 2017), PNN (Rusu et al. 2016), FMR (Zhu et al. 2024), as well as energy-based and fine-grained token-driven models ETT-CKGE (Zhu et al. 2025) and CLKGE (Cao 2024).
@@
Metrics We evaluate link prediction performance using MRR, Hits@1, Hits@3, and Hits@10, where higher is better. Following (Bordes et al. 2013) for negative sampling, we rank candidate triples by replacing the head or tail entity. For each snapshot $i$, we average scores over the current and all previous test sets. Results are reported using the fully trained model.
Results and Analysis

### v2

@@ 4.2. Bayesian Guided Knowledge Evolution Learning
Inspired by related work such as uncertainty regularized continual learning(Ahn et al., 2019)(Lee et al., 2024b)(Bonnet et al., 2025), we regard the embedding parameters of entities and relations as random variables in probability distributions rather than fixed point estimates. This approach not only facilitates quantification of parameter uncertainty but also protects important old knowledge when learning new knowledge, avoiding catastrophic forgetting.
@@
Entity and Relation Distributional Representations. For each entity $e$ and relation $r$, their embedding vectors at time snapshot $t$ are modeled as Gaussian distributions. The embedding of entity $e$ is represented as $\mathbf{e}_{t}\sim\mathcal{N}(\boldsymbol{\mu}_{e,t},(\text{diag}(\boldsymbol{\lambda}_{e,t}))^{-1})$, where $\boldsymbol{\mu}_{e,t}\in\mathbb{R}^{d}$ is the mean vector, and $\boldsymbol{\lambda}_{e,t}\in\mathbb{R}^{d}$ is the diagonal precision vector (inverse of variance); the embedding of relation $r$ is similar, represented as $\mathbf{r}_{t}\sim\mathcal{N}(\boldsymbol{\mu}_{r,t},(\text{diag}(\boldsymbol{\lambda}_{r,t}))^{-1})$. Each element of the precision $\boldsymbol{\lambda}$ reflects the certainty of the corresponding dimension: a higher value indicates more reliable knowledge in that dimension, which should be better protected in subsequent updates.
@@
Sequential Bayesian Update. At time snapshot $t$, we use the posterior distribution $\{\boldsymbol{\mu}_{t-1},\boldsymbol{\lambda}_{t-1}\}$ from time t1t-1 as the prior and perform online Bayesian update using new data $\Delta\mathcal{T}_{t}$. To simplify inference, the embeddings $\hat{\boldsymbol{\Theta}}_{t}$ obtained from training on the current snapshot are treated as new observations to the prior mean $\boldsymbol{\mu}_{t-1}$, generated by the classic KGE model TransE (Bordes et al., 2013). TransE evaluates the plausibility of triples through a scoring function and optimizes the embeddings $\hat{\boldsymbol{\Theta}}_{t}=\{\hat{\mathbf{e}}_{t},\,\hat{\mathbf{r}}_{t}\}$ by minimizing the margin-based ranking loss $\mathcal{L}_{\text{KGE}}$:
(1) $\small\mathcal{L}_{\mathrm{KGE}}=\sum_{(h,r,t)\in\mathcal{T}}\sum_{\left(h^{\prime},r,t^{\prime}\right)\in\mathcal{T}^{-}}\left[\gamma+f(h,r,t)-f\left(h^{\prime},r,t^{\prime}\right)\right],$
@@
where $f(h,r,t)=\lVert\mathbf{h}+\mathbf{r}-\mathbf{t}\rVert_{2}$ represents the score function of TransE(Bordes et al., 2013). Based on the conjugate property of Gaussian distributions, the update rules for the embedding parameters of entity $e$ are:
(2) $\boldsymbol{\lambda}_{e,t}$ $=\boldsymbol{\lambda}_{e,t-1}+\lambda_{obs},$
(3) $\boldsymbol{\mu}_{e,t}$ $=\frac{\boldsymbol{\lambda}_{e,t-1}\odot\boldsymbol{\mu}_{e,t-1}+\lambda_{obs}\odot\hat{\mathbf{e}}_{t}}{\boldsymbol{\lambda}_{e,t}},$
@@
To guide the embeddings towards the posterior mean, we use a regularization term (Lee et al., 2024b) $\mathcal{L}_{Bayes}$, which is equivalent to minimizing the KL divergence between the current distribution $q(\mathbf{\Theta}_{t})$ and the target posterior:
(4) $\mathcal{L}_{Bayes}=\sum_{i\in\mathcal{E}_{t}\cup\mathcal{R}_{t}}\beta\cdot\left\|\sqrt{\boldsymbol{\lambda}_{i,t-1}}\odot(\hat{\boldsymbol{\theta}}_{i,t}-\boldsymbol{\mu}_{i,t-1})\right\|_{2}^{2},$
@@
4.3. Continual Clustering
@@
Bayesian updates prevent forgetting at the parameter level, but the evolution of knowledge representations between KG snapshots needs to be constrained; otherwise, entity and relation embeddings in the latent space may undergo uncertainty drift across snapshots. Inspired by contrastive learning and clustering methods (Liang et al., 2024a)(Truong et al., 2025)(Nguyen et al., 2021), we propose a continual clustering method as a constraint to maintain the geometric structural consistency of the embedding space at the semantic level.
@@
Sequential Contrastive Clustering Learning. In continual learning, imbalances in data volume and class distribution across temporal snapshots may lead to under-learning of minority classes. To address this, we propose the sequential contrastive clustering method ($\mathcal{L}_{FCC}$), which achieves fairness by dynamically adjusting the contributions of major and minor classes. Specifically, for each cluster $k$, we maintain a centroid vector $\mathbf{c}_{k}\in\mathbb{R}^{d}$ (where $d$ is the embedding dimension) to represent the central features of the cluster. First, we acknowledge the perspective proposed in (Liu et al., 2024a) that the learning order is also crucial in the process of learning KG structures. Second, since new entities emerge in each snapshot, making it difficult to apply k-means clustering, we sort the entities based on their importance, then fix the size of each cluster $k$, and take the mean of the embeddings in the current cluster $k$ as the clustering centroid $\mathbf{c}_{k}$. Specifically, the sorting process is as follows: we comprehensively consider the node’s centrality and betweenness centrality in the graph structure to compute its importance score. The calculation formulas for these two metrics are as follows:
(5) $f_{nc}(e)=\frac{f_{\text{neighbor}}(e)}{N-1},\quad f_{bc}(e)=\sum_{s,t\in\mathcal{E},s\neq t}\frac{\sigma(s,t|e)}{\sigma(s,t)}$
@@
where $f_{nc}(e)$ reflects the tightness of local connections for the entity, and $f_{bc}(e)$ measures its bridging role in information propagation. The variables in the formulas are defined as follows: $f_{\text{neighbor}}(e)$ represents the number of neighbors of entity $e$, NN is the total number of entities in the current knowledge graph snapshot, $\mathcal{E}$ is the set of entities in the current time snapshot $i$, $\sigma(s,t)$ is the total number of shortest paths between entities $s$ and $t$, and $\sigma(s,t\mid e)$ is the number of those paths that pass through entity $e$. We combine the two to define the importance score for entity $e$ as: $IE(e)=f_{nc}(e)+f_{bc}(e)$. We calculate the IE score for each entity and sort them in descending order of scores. Then, we assign the sorted entities to clusters of fixed size in sequence, and compute the initial embedding for each cluster as $\mathbf{c}_{k}=\frac{1}{|K|}\sum_{\mathbf{e}_{i}\in K}\mathbf{e}_{i}$. The continual clustering then constrains the representations of entities across different snapshots through contrastive loss, with the loss function defined as:
@@
(6) $\small\mathcal{L}_{\text{FCC}}=-\sum_{k=1}^{K}\Bigl(\alpha_{k}\!\!\sum_{\mathbf{e}_{i}\in\text{Cluster}_{k}}\mathcal{L}_{\text{cont}}(\mathbf{e}_{i},\mathbf{c}_{k})+\mathcal{L}_{\text{cont}}(\mathbf{v}_{k},\mathbf{c}_{k})\Bigr)$
where $\mathbf{e}_{i}$ is the entity embedding vector belonging to cluster $k$, $\text{sim}(\cdot,\cdot)$ is the cosine similarity function, $\tau$ is the temperature parameter (controlling the smoothness of softmax), $K$ is the total number of clusters (equal to the number of entity classes in the current snapshot), and $\mathcal{L}_{\text{cont}}(\mathbf{e}_{i},\mathbf{c}_{k})$ is the standard contrastive loss, defined as:
@@
(7) $\mathcal{L}_{\text{cont}}\left(\mathbf{e}_{i},\mathbf{c}_{k}\right)=-\log\frac{\exp\left(\operatorname{sim}\left(\mathbf{e}_{i},\mathbf{c}_{k}\right)/\tau\right)}{\sum_{j=1}^{K}\exp\left(\operatorname{sim}\left(\mathbf{e}_{i},\mathbf{c}_{j}\right)/\tau\right)}.$
@@
Cluster Maintenance and Update. At snapshot $t$, the centroids $\mathbf{c}_{k}$ of old classes are inherited and fixed from t1t-1, serving as anchors for old knowledge. New classes initialize centroids via feature means. During training, entities are dynamically assigned to the nearest cluster, and centroids are slowly adjusted through a momentum update:
(8) $\mathbf{c}_{k,t}=(1-\eta)\mathbf{c}_{k,t-1}+\eta\cdot\text{mean}(\{\mathbf{e}_{i}\in\text{Cluster}_{k}\}),$
@@
4.4. Final Training Objective
@@
(9) $\mathcal{L}_{total}=\mathcal{L}_{KGE}+\mathcal{L}_{Bayes}+\mathcal{L}_{FCC}.$
5. Experiments
5.1. Experimental Setup
Datasets. To thoroughly verify the effectiveness and stability of BAKE, we conducted extensive experiments on a wide variety of datasets, including: ENTITY (Cui et al., 2023), RELATION (Cui et al., 2023), HYBRID (Cui et al., 2023), GraphLower (Liu et al., 2024a), GraphEqual (Liu et al., 2024a), GraphHigher (Liu et al., 2024a), FB-CKGE (Liu et al., 2024b), and WN-CKGE (Liu et al., 2024b). These eight datasets encompass different evolution patterns, each with unique characteristics. Detailed statistics of the datasets are listed in TABLE 5. Each dataset contains 5 snapshots.
Implementation Details. All experiments are implemented in PyTorch(Paszke et al., 2019) and conducted on 8 NVIDIA Tesla V100 GPUs. We tune all hyperparameters via grid search, including embedding dimensions for entities and relations $D\in\{50,100,200,300\}$, batch sizes $B\in\{256,512,1024\}$, and number of clusters $K$ $\in\{256,512,1024\}$, among others. We use Adam as the optimizer with learning rates selected from {1e-5, 1e-4, 1e-3}. To ensure fairness, all reported results are averaged over five runs.
Table 1. Performance comparison of various models on ENTITY, RELATION, FB-CKGE, and WN-CKGE datasets. The best results are in bold, and the second-best results are underlined.
@@ 5.1. Experimental Setup
ETT-CKGE 0.260 0.158 - 0.456 0.200 0.112 - 0.369 0.236 0.137 - 0.428 0.153 0.008 - 0.385
CLKGE 0.248 0.144 0.278 0.436 0.203 0.115 0.226 0.379 - - - - - - - -
SAGE 0.280 0.176 - 0.477 0.217 0.122 - 0.397 - - - - - - - -
BAKE 0.287 0.179 0.326 0.484 0.224 0.126 0.239 0.387 0.257 0.157 0.299 0.451 0.172 0.019 0.295 0.404
Improve 2.5% 1.7% 12.0% 1.5% 3.2% 3.3% 5.8% - 8.9% 14.6% 10.3% 5.4% 8.2% 26.7% 2.8% 4.9%
Table 2. Performance comparison of various models on GraphLower, GraphEqual, GraphHigher, and HYBRID datasets.
@@
SAGE 0.237 0.138 0.432 0.247 0.147 0.446 0.239 0.143 0.426 0.224 0.131 0.400
BAKE 0.247 0.147 0.448 0.257 0.154 0.463 0.241 0.145 0.439 0.228 0.146 0.419
Improve 4.2% 6.5% 3.7% 4.0% 4.8% 3.8% 0.8% 1.4% 3.0% 1.8% 11.5% 4.5%
Table 3. Ablation study results on ENTITY, RELATION, FB-CKGE, and WN-CKGE datasets.
ENTITY RELATION FB-CKGE WN-CKGE
@@
Baselines. To fully demonstrate the effectiveness of BAKE, we compare it with all recent CKGE models, including Fin-tune (Cui et al., 2023) (based on fine-tuning with new triples), IncDE (Liu et al., 2024a) and FastKGE (Liu et al., 2024b) (based on incremental distillation and LoRA fine-tuning), LKGE (Cui et al., 2023), EWC (Kirkpatrick et al., 2017), SI (Zenke et al., 2017), DiCGRL (Kou et al., 2020), EMR (Wang et al., 2019), CWR (Lopez-Paz and Ranzato, 2017), PNN (Rusu et al., 2016), FMR (Zhu et al., 2024), as well as energy-based and fine-grained token-driven models ETT-CKGE (Zhu et al., 2025), SAGE(Li et al., 2025b) and CLKGE (Cao, 2024).
@@
Metrics. We evaluate link prediction performance using MRR, Hits@1, Hits@3, and Hits@10, where higher is better. Following (Bordes et al., 2013) for negative sampling, we rank candidate triples by replacing the head or tail entity. For each snapshot $i$, we average scores over the current and all previous test sets. Results are reported using the fully trained model.
5.2. Results and Analysis